OpenStudy (anonymous):

what is the slope of these two equations. 4x+5y=15 16x+20y=0

OpenStudy (nikato):

the easiest way is to change the equations to slope intercept form whcih is] y=mx+b where m=slope i'll do the first one 4x+5y=15 -4x -4x -------------- 5y=-4x+15 divide both sides by 5 y=-4/5x +3 so now its in y=mx+b form and your m=-4/5 so your slope is -4/5 can you try to do the other one?

OpenStudy (anonymous):

-4/5 qagain